Development Meeting

(Seminar Room, Buccleuch Place)- Minutes 2008-04-02

Present: Toby Blake, Roger Burroughes, Chris Cooke, Ken Dawson (convener and minutes), Ian Durkacz, Iain Rae, George Ross, Craig Strachan

Apologies: Alastair, Neil, Simon, Stephen and Tim had sent their apologies.

1. Minutes of last meeting

These were accepted.

2. Matters Arising

Seventeen out of the twenty seven computing staff had now moved their mail service to staffmail.

Paul has now moved but Simon will not be able to move for another month or so. Neil reported that the Default quota had now been increased to 2 GB. However this was a limit that could not be exceeded until IS introduce a mail server with a later version of Cyrus software.

Richard and John have also still to move.

Richard was going to try to move today.

Craig commented that one of the non-computing staff on the Computing Strategy Group had now also moved to staffmail. There was no longer a 2GB limit to mail boxes and new staff will only get a staffmail mail account; an alias for their inf mail address pointing at their staffmail address will automatically be inserted into the alias file for the inf mail server.

Simon had still to migrate to staffmail. [Action Simon]

Craig will find out if there is any substantive reason why Simon has not moved. [Action Craig]

  • Revisit Account Management Simon had produced the discussion paper and the next step will be for George to call a meeting of interested parties. The subject will be discussed at the end of this meeting and the need for and nature of any further meeting will be dependent on the outcome of this.

Because of Simon's absence it had not been possible for George to speak to him about this.

Simon pointed out that the meeting was needed to sort out the short term problem of instituting a procedure to remove any extraordinary kerberos principles belonging to people who have left. George will arrange a meeting between himself, Simon and Ken.

George is yet to arrange this. [Action George]

* DICE SL5 Server Platform There were two pieces of software that had not apparently been ported to the SL5 64bit platform (the SL5 platform was complete). Toby confessed that he had actually ported Cosign to SL5-64 but had not closed down the corresponding bugzilla bug. Stephen offered to look at the outstanding problem. Ian will talk to Stephen about this.

George has still to port his tnm package to 64-bit. This software is in the tkined rpm. Ian has been looking into compiling this on a 64-bit SL5 platform. [Action George]

  • Cluster Parallel filesystem Iain reported that an issue had emerged about licensing of the software and that this needed to be clarified. He will discuss with Tim and Alastair.

Iain has spoken to Alastair about this and both have written to IBM.

All other actions from the last meeting have been completed.

3. Approved Projects

  • AFS file system The next milestone 'All Staff on AF' has not been achieved yet because of the pressure on User Support staff. Approximately 200 staff have been moved to AFS but there are still about 300 still using NFS. Craig has discussed with Alison whether we should continue with moving staff one by one or do a mass move. The latter would almost certainly place even more pressure on the User Support staff and so it was agreed to continue moving staff on an individual basis for the foreseeable future. We should revisit this issue in August. Craig should reset the milestone proposed date for 'All Staff on AF' to sometime in August.

  • Web Technologies and Policy This project has effectively been completed. Craig will submit it for sign off for the next meeting. A new project will be proposed to explore how we move forward with web technologies in the future.

  • Infrastructure for New Informatics Forum The deadline for handing over the Forum to the University has passed but only parts of floors 2 and 3 have been handed over so far. The technicians have finished wiring the three IT closets on floor 2 and have started work on the IT closets on floor 3. Work has begun on configuring the switches for floor 4. All switches for the Forum have now been delivered.

  • DICE SL5 Services The Managed Platform Unit and the Infrastructure Unit have upgraded all their servers. The other three units envisage upgrading their remaining servers in the next couple of weeks (although the situation with the Services Unit's Linux NFS servers is still a little uncertain).

4. New project proposals

  • SL5 Desktop The project proposal was accepted and the project moves forward to the implementation stage.

5. AOB

The next meeting will be on May 7th in the Room 2511, JCMB.

-- KenDawson - 02 Apr 2008

Topic revision: r2 - 04 Apr 2008 - 09:29:42 - KenDawson
 
This site is powered by the TWiki collaboration platformCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback
This Wiki uses Cookies